# Learning Spatial-Semantic Features for Robust Video Object Segmentation

### v1

@@ 5 Limitations
Our method enhances target representation by introducing spatial-semantic features and discriminative query propagation, achieving significant improvements in numerous challenging scenarios. However, in the scenes with object parts as the target (some sequences in the VOT challenge, shown in the appendix), the proposed model finds it difficult to represent a part of the target and often extends the predicted mask to the whole target. For example, when the target is the head of a person, our model may gradually shift to segmenting the entire body. This tendency arises because our method tends to interpret the target as the entire object seen during the training phase. A potential direction to improve this issue could be letting the model learn more part-aware semantic features, which also needs a new benchmark.
@@ 6 Conclusion
In this paper, we propose a spatial-semantic network to learn comprehensive representations to improve the robustness of video object segmentation in handling objects with complex structures or separate parts. In addition, we develop a discriminative query propagation module that selects representative features in target objects for query update, enabling more reliable query propagation in long-term videos. Our approach achieves favorable performance compared with the state-of-the-art methods on numerous datasets. Both quantitative and qualitative results demonstrate the effectiveness of the proposed method in tracking and segmenting video objects in long-term and complex scenes.
@@ A.1.1 Multi-Object Processing
Following the previous work [10, 11], we extend our methods to the multi-object setting. The predicted mask of each frame undergoes channel separation for different targets. The separated masks for each target are then concatenated with the RGB image and the aggregated mask of other targets for feature extraction. The extracted features are fused with RGB features to obtain pixel-level value features, which are then stored in memory. It’s important to note that the pixel-level features stored in memory include both the key (RGB feature) and the value.
@@ A.1.2 Training Details
As mentioned in the main paper, we train our network only on video-level datasets, which is different from other methods [14, 10, 51]. The backbone weights are initialized from MAE pretraining, similar to the approaches used in SimVOS [44] and JointFormer [32]. We conduct spatial-semantic fusion every three ViT layers and five discriminative query transformer blocks for query propagation. We implement our network using PyTorch and train the model on 8 NVIDIA V100s, employing automatic mixed precision (AMP) during training.
Two different training datasets are used for better improvements. The "DAVIS&\&&YTVOS" settings combine the training sets of YouTubeVOS-2019 and DAVIS-2017, with the DAVIS-2017 dataset expanded five times to increase the data ratio. The "MEGA" set contains four different training sets, including DAVIS-2017, YouYubeVOS-2019, MOSE, and BURST datasets, with the DAVIS-2017 dataset expanded five times to increase the data ratio. To sample a training sequence, we first randomly select a "seed" frame from all the frames and then randomly select 7 other frames from the same video. We re-sample if any two consecutive frames have a temporal distance greater than the max-skip. Following the previous work [10, 14], the max-skip is set to [5, 10, 15, 5] after [0%, 10%, 30%, 80%] of the training iterations, respectively. To avoid sampling frames without targets, we use the data statistics provided by Cutie [14].
Table A1: Training parameters.
@@
In the training phase, we initialize the model with the first frame and ground truth. When the number of past frames is three or fewer, we use all of them as memory frames. If there are more than three past frames, we randomly sample three of them to be the memory frames. We compute the loss at all frames except the first one and perform back-propagation through time.
Table A2: Additional ablation study.
@@
XMem [10] (Baseline) 53.3 62.0 57.6 81.0 77.4 84.5 50.0 45.5 54.4 85.5 84.3 88.6 80.3 88.6
Cutie [14] 64.0 60.0 67.9 84.2 80.6 87.7 56.2 51.8 60.5 86.1 85.8 90.5 80.0 88.0
@@
A.1.3 Inference Details
During the inference phase, the model memorizes the feature every 5th frame like  [10, 14]. The pixel-level memory always stores the first frame and its mask. The top-K filter is used to augment the memory-reading in pixel memory. For object-level memory, we use the streaming average to accumulate the discriminative query representation like the operation in  [14].
A.2 Additional Quantitative Results
In this section, we provide more additional ablation studies and sota comparisons with different settings. In the main paper, we provide ablation studies for different components. To provide a broader comparison, we conduct tests under various combinations of training data and testing settings. The experimental results in Table A2 detail the impact of different components of our model on the performance, facilitating a better understanding of the effectiveness of our model.
To compare with more methods on MOSE, we follow the training strategy from MOSE [17], and replace the YouTubeVOS with MOSE in the training process. Table A3 shows the results of the variants of our method. All the variants achieve great improvement in terms of $\mathcal{J}\&\mathcal{F}$, which further validates the effectiveness of the proposed spatial-semantic feature learning and discriminative query representation.
Table A3: Results only trained on MOSE datasets.
@@ A.2 Additional Quantitative Results
RDE [27] 48.8 44.6 52.9
STCN [11] 50.8 46.6 55.0
AOT [50] 57.2 53.1 61.3
XMem [10] 57.6 53.3 62.0
DeAOT [51] 59.4 55.1 63.8
@@
Figure A.1: Additional qualitative comparison against the state-of-the-art methods including XMem, Swin-DeAOT, and Cutie. All the sequences are picked from YouTubeVOS 2019 and DAVIS 2017 datasets.
Figure A.2: Additional qualitative comparison on sequence with giant appearance variance. All the sequences are picked from MOSE [17] dataset.
@@ A.3 Additional Qualitative Comparison
We present a more qualitative comparison of most benchmarks in Figure A.1 and Figure A.2.
@@
In Figure A.1, we categorize the scenarios into three main types: overall-to-part segmentation (man on bike and deer), small object segmentation (small monkey), blurry object segmentation (cat toy), and segmentation under varying lighting conditions (man-bike and golf). In these three cases, our method demonstrates precise segmentation compared to other methods. The semantic embedding allows our model to consider semantic information when parsing objects, spatial dependence modeling enhances the understanding of spatial positions and details of the object, and the discriminative query accurately represents the target. The integration of these three components enables the model to achieve excellent segmentation and tracking performance in complex and challenging scenarios.
@@
In Figure A.2, we select more sequence results on MOSE, primarily divided into two categories: positional changes of similar targets and rapid movement of targets. In both cases, our method outperforms Cutie. This is attributed to our proposed spatial-semantic feature representation and discriminative query expression, which enable the model to distinguish appearance-similar targets and consistently track the target.
@@
Figure A.3 compares sequences that require a semantic understanding of the target. Without the semantic embedding module, the method suffers from segmenting parts of the target. The addition of the semantic embedding module enables the model to fully understand the semantic information of the target, allowing for precise segmentation and tracking in these scenarios.
@@
Failure cases are shown in Figure A.4, in which our method tends to segment the whole instance rather than the part of them. The challenging sequences need the model to segment and track the head of the man, flamingo, and kangaroo. Our model interprets semantics as the entire instance during tracking and segmentation, such as a person, bird, or kangaroo, resulting in segmentation outcomes gradually favoring the whole instance.
@@
Figure A.3: Qualitative comparison between with and without semantic embedding block. All the sequences are picked from YouTubeVOS 2019 dataset.
@@
Figure A.4: Failure cases. All the sequences are picked from VOT Challenge [26].

### v2

@@ 4.4 Limitations
Our method enhances target representation by introducing spatial-semantic features and discriminative query propagation, significantly improving numerous challenging scenarios. However, in scenes with object parts as the target (some sequences in the VOT challenge, shown in the appendix), the proposed model finds it difficult to represent a part of the target and often extends the predicted mask to the whole target. For example, when the target is the head of a person, our model may gradually shift to segmenting the entire body. This tendency arises because our method tends to interpret the target as the entire object seen during the training phase. A potential direction to improve this issue could be to let the model learn more part-aware semantic features, which also needs a new benchmark. Moreover, the improvement of the inference speed will also be addressed in our future work.
@@ 5 Conclusion
In this paper, we propose a spatial-semantic network to learn comprehensive representations to improve the robustness of video object segmentation in handling objects with complex structures or separate parts. In addition, we develop a discriminative query propagation module that selects representative features from target objects for query update, enabling more reliable query propagation in long-term videos. Our approach achieves favorable performance compared with state-of-the-art methods on numerous datasets. Both quantitative and qualitative results demonstrate the effectiveness of the proposed method in tracking and segmenting video objects in complex and long-term scenes.
Acknowledgments This work is supported by the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515011292), the National Natural Science Foundation of China (Grant No.62476148, 62172126), and the Shenzhen Research Council (No. JCYJ20210324120202006).
@@ A.1.1 Multi-Object Processing
Following the previous work (Cheng & Schwing, 2022; Cheng et al., 2021a), we extend our methods to the multi-object setting. The predicted mask of each frame undergoes channel separation for different targets. The separated masks for each target are then concatenated with the RGB image and the aggregated mask of other targets for feature extraction. The extracted features are fused with RGB features to obtain pixel-level value features, which are then stored in memory. It is important to note that the pixel-level features stored in memory include the key (RGB feature) and the value.
@@ A.1.2 Training Details
As mentioned in the main paper, we train our network only on video-level datasets, which differs from other methods (Cheng et al., 2023a; Cheng & Schwing, 2022; Yang & Yang, 2022). The backbone weights are initialized from MAE pre-training, similar to the approaches used in SimVOS (Wu et al., 2023) and JointFormer (Mao et al., 2021). We perform spatial-semantic fusion every three ViT layers and five discriminative query transformer blocks for query propagation. We implement our network using PyTorch and train the model on 8 NVIDIA V100s, employing automatic mixed precision (AMP) during training.
Two different training datasets are used for better improvements. The "DAVIS&\&&YTVOS" settings combine the training sets of YouTubeVOS-2019 and DAVIS-2017, with the DAVIS-2017 dataset expanded five times to increase the data ratio. The "MEGA" set contains four different training sets, including DAVIS-2017, YouYubeVOS-2019, MOSE, and BURST datasets, with the DAVIS-2017 dataset expanded five times to increase the data ratio. To sample a training sequence, we randomly select a "seed" frame from all the frames and then randomly choose 7 other frames from the same video. We re-sample if any two consecutive frames have a temporal distance greater than the max-skip. Following the previous work (Cheng & Schwing, 2022; Cheng et al., 2023a), the max-skip is set to [5, 10, 15, 5] after [0%, 10%, 30%, 80%] of the training iterations, respectively. To avoid sampling frames without targets, we use the data statistics provided by Cutie (Cheng et al., 2023a).
Table 5: Training parameters.
@@
We initialize the model with the first frame and ground truth in the training phase. When the number of past frames is three or fewer, we use all of them as memory frames. If there are more than three past frames, we randomly sample three of them to be the memory frames. We compute the loss at all frames except the first one and perform back-propagation through time.
Table 6: Quantitative comparisons with high-resolution input on the MOSE, LVOS, DAVIS 2017, YouTube-VOS 2018 & 2019 dataset. + means improve the input size.
Dataset MOSE-val DAVIS 2017 test YouTube-VOS 2018 val YouTube-VOS 2019 val
[…]
Cutie-base (Cheng et al., 2023a)* 64.0 60.0 67.9 84.2 80.6 87.7 86.1 85.8 90.5 80.0 88.0 86.1 85.5 90.0 80.6 88.3
ISVOS (Wang et al., 2023)*+BL30K (Cheng et al., 2021b) - - - 84.0 80.1 87.8 86.7 86.1 90.8 81.0 89.0 86.3 85.2 89.7 81.0 89.1
JointFormer (Zhang et al., 2023)*+BL30K (Cheng et al., 2021b) - - - 88.1 84.7 91.6 87.6 86.4 91.0 82.2 90.7 87.4 86.5 90.9 82.0 90.3
Ours 68.5 64.5 72.6 87.1 83.1 91.1 87.4 87.0 92.0 80.9 89.7 87.5 86.8 91.8 81.3 89.9
Cutie-base (Cheng et al., 2023a)*+ 66.2 62.3 70.1 85.9 82.6 89.2 - - - - - 86.9 86.2 90.7 81.6 89.2
Ours+ 70.5 66.5 74.6 87.9 84.6 91.3 87.6 86.9 91.7 81.5 90.1 87.8 86.8 91.6 82.2 90.8
Cutie-base* (Cheng et al., 2023a) w/ MEGA 69.9 65.8 74.1 86.1 82.4 89.9 87.0 86.4 91.1 81.4 89.2 87.0 86.0 90.5 82.0 89.6
Ours w/MEGA 73.2 68.8 77.5 88.2 84.3 92.1 88.1 87.4 92.5 81.9 90.7 88.0 88.0 91.8 82.5 90.8
Cutie-base (Cheng et al., 2023a)*+ w/ MEGA 71.7 67.6 75.8 88.1 84.7 91.4 - - - - - 87.5 86.3 90.6 82.7 90.5
Ours+ w/MEGA 75.1 71.0 79.2 89.1 85.8 92.4 88.5 87.6 92.6 82.7 91.3 88.5 87.3 92.0 83.1 91.4
A.1.3 Inference Details
During the inference phase, the model memorizes the feature every 5th frame like  (Cheng & Schwing, 2022; Cheng et al., 2023a). The pixel-level memory always stores the first frame and its mask. The top-K filter is used to augment memory reading in pixel memory. We use the streaming average for object-level memory to accumulate the discriminative query representation like the operation in  (Cheng et al., 2023a).
A.2 Additional Quantitative Results
In this section, we provide additional ablation studies and SOTA comparisons with different settings.
Table 6 shows detailed results about methods with further pre-training (e.g. JointFormer (Zhang et al., 2023) and ISVOS (Wang et al., 2023)) and improved test size (Cutie (Cheng et al., 2023a))We note that our method achieves comparable results against SOTA methods (Cutie (Cheng et al., 2023a) and JointFormer (Zhang et al., 2023)) without any pre-training. When trained with MEGA datasets or improved test image size, our method achieves state-of-the-art results in all the benchmarks.
In the main paper, we provide ablation studies for different components. To provide a broader comparison, we conduct tests using various combinations of training data and testing settings. The experimental results in Table 7 detail the impact of different components of our model on the performance, facilitating a better understanding of the effectiveness of our model.
Table 7: Model detailed ablation study.
@@ A.2 Additional Quantitative Results
XMem (Cheng & Schwing, 2022) (Baseline) 53.3 62.0 57.6 81.0 77.4 84.5 50.0 45.5 54.4 85.5 84.3 88.6 80.3 88.6
Cutie (Cheng et al., 2023a) 64.0 60.0 67.9 84.2 80.6 87.7 56.2 51.8 60.5 86.1 85.8 90.5 80.0 88.0
@@
To compare with more methods on MOSE, we follow the training strategy from MOSE (Ding et al., 2023), and replace the YouTubeVOS with MOSE in the training process. Table 8 shows the results of the variants of our method. All the variants achieve great improvement in terms of $\mathcal{J}\&\mathcal{F}$, which further validates the effectiveness of the proposed spatial-semantic feature learning and discriminative query representation.
Table 8: Results only trained on MOSE datasets.
@@
RDE (Li et al., 2022a) 48.8 44.6 52.9
STCN (Cheng et al., 2021a) 50.8 46.6 55.0
AOT (Yang et al., 2021) 57.2 53.1 61.3
XMem (Cheng & Schwing, 2022) 57.6 53.3 62.0
DeAOT (Yang & Yang, 2022) 59.4 55.1 63.8
@@
Table 9 shows the results of our method on the more challenging VOTS2023 (Kristan et al., 2023) benchmark, which indicates that our method outperforms the winner (DMAOT) of last year on most metrics. Our method achieves the highest robustness thanks to our representation of target saliency information and the modeling of target semantic-spatial information.
Table 9: Results on VOTS2023 Challenge.
Dataset VOTS2023
Method $Q\uparrow$ $A\uparrow$ $R\uparrow$ $NER\downarrow$ $DRE\downarrow$ $ADQ\downarrow$
Dynamic_DEAOT 0.59 0.69 0.84 0.07 0.1 0.57
M-VOSTrack 0.61 0.75 0.76 0.16 0.08 0.71
HQTrack 0.62 0.75 0.77 0.16 0.08 0.70
DMAOT 0.64 0.75 0.80 0.14 0.07 0.73
Ours 0.67 (+0.3) 0.74 0.86 (+0.6) 0.09 (+0.5) 0.05 (+0.02) 0.75 (+0.02)
Figure 5: Additional qualitative comparison against state-of-the-art methods including XMem, Swin-DeAOT, and Cutie. All the sequences are selected from YouTubeVOS 2019 and DAVIS 2017 datasets.
Figure 6: Additional qualitative comparison on sequence with giant appearance variance. All the sequences are picked from MOSE (Ding et al., 2023) dataset.
@@ A.3 Additional Qualitative Comparison
We present a more qualitative comparison of most benchmarks in Figure 5 and Figure 6.
@@
In Figure 5, we categorize the scenarios into three main types: overall-to-part segmentation (man on bike and deer), small object segmentation (small monkey), blurry object segmentation (cat toy), and segmentation under varying lighting conditions (man-bike and golf). In these three cases, our method demonstrates precise segmentation compared to other methods. The semantic embedding allows our model to consider semantic information when parsing objects, spatial dependence modeling enhances the understanding of spatial positions and details of the object, and the discriminative query accurately represents the target. The integration of these three components enables the model to achieve excellent segmentation and tracking performance in complex and challenging scenarios.
@@
In Figure 6, we select more sequence results on MOSE, primarily divided into two categories: positional changes of similar targets and rapid movement of targets. In both cases, our method outperforms Cutie. This is attributed to our proposed spatial-semantic feature representation and discriminative query expression, which enable the model to distinguish appearance-similar targets and consistently track the target.
@@
Figure 9 compares sequences that require a semantic understanding of the target. Without the semantic embedding module, the method suffers from segmenting parts of the target. The addition of the semantic embedding module enables the model to fully understand the semantic information of the target, allowing for precise segmentation and tracking in these scenarios.
Visualization of feature maps. Figure 7(a) shows that the proposed Spatial-Semantic block assigns higher activation weights to complex target objects compared to the CNN or ViT model, which indicates that the proposed model makes target representations more distinguishable than ViT and ResNet. Moreover, we show how the feature changes during the processing of the proposed model by visualizing the feature maps at different processing steps in  7(b). After passing through the semantic embedding module, the feature maps (semantic embedding) demonstrate that the module makes the contours of objects clearer, indicating richer semantic information. After passing through the spatial dependence module, the feature maps (spatial modeling) show that the details of the targets are more pronounced. The changes in the feature maps confirm that the proposed two modules significantly enhance the semantic and spatial information of the targets.
We also visualize the similarity between the learned discriminative queries and the feature map using heat maps in Figure 8. The visualization shows that the learned discriminative queries distinctly express the target areas. We also mark the positions of the discriminative queries, which represent the locations of the pixels with the highest similarity.
@@
Failure cases are shown in Figure 10, in which our method tends to segment the whole instance rather than the part of it. The challenging sequences require the model to segment and track the head of the man, the flamingo, and the kangaroo. Our model interprets semantics as the entire instance during tracking and segmentation, such as a person, bird, or kangaroo, resulting in segmentation outcomes gradually favoring the whole instance.
(a) Visualization of the feature map output from different backbones
(b) Visualization of the feature map after proposed semantic and spatial layers
Figure 7: Visualizations of feature maps from different architectural components.
Figure 8: Visualization of the similarity map between query and features. The map indicates the discrimination of the target representation. The discriminative queries of different targets are plotted in the similarity map, which can be found in the black box.
@@
Figure 9: Qualitative comparison between with and without semantic embedding block. All the sequences are picked from the YouTubeVOS 2019 dataset.
@@
Figure 10: Failure cases. All the sequences are picked from VOT Challenge (Kristan et al., 2023).